This is a Mixed question / It is worth 1 point / You have 4 of 4 attempts remaining / There is a 15% attempt penalty

See page 398

07Question(1 point)

1st attempt

Part 1   (0.5 point)

See Hint

See Periodic Table

Which of the following forms the most polar bond?

blank.pngChoose one: B and Cl  F and C Cl and O H and F

Part 2   (0.5 point)

See Hint

Are any of the bonds formed between the following pairs of atoms ionic?

blank.pngChoose one or more:

O and H

Ca and Cl

P and F

B and H

C and Cl
Answer #1

81 B-ce electronegativity difference AX = 3-2 = 1 C-F at DX = 4-2.5 = 1.5 Cl-o o x = 3.5-3.0 = 0.5 H-F LAX = 4.0-2.1 - 1.9 @nCa-Cl and P-F bond are ionic.